# ACDC: The Adverse Conditions Dataset with Correspondences for Robust Semantic Driving Scene Perception

### v4

@@ 3.5 Instance Annotation
The total 4006 adverse-condition images and 1503 normal-condition reference images are annotated at the instance level by a professional team of annotators. Each final annotation is checked by at least two annotators and passed through quality control. The total time required for annotating a single image given the initial semantic mask was 0.6 h on average.
@@ 3.6 Comparison to Related Datasets
To the best of our knowledge, ACDC constitutes the largest adverse-condition driving dataset for dense semantic perception to date. In Table I, we compare ACDC to prior datasets that also address semantic segmentation under adverse conditions. Most of these datasets focus on a single condition and are of small scale. WildDash covers a wider variety of adverse conditions but also has a small scale. BDD100K includes 10000 images with semantic segmentation annotations. We inspected these images manually to identify those that pertain to fog, night, rain, and snow. We found that only 1346/10000 images pertain to any of these four conditions. By contrast, ACDC is primarily composed of these four common adverse conditions. Notably, it contains one order of magnitude more annotated images than any other competing dataset for each of fog, night and rain. At the same time, our specialized annotation protocol using corresponding normal-condition images ensures reliable annotations even for invalid regions, making ACDC a high-quality dataset for training and evaluation for adverse conditions. We also provide a comparison on the number of annotated instances with existing datasets covering adverse conditions in Table II. Only a small portion of BDD100K provides instance-level annotations in adverse conditions including fog, night, rain and snow. The DAWN dataset consists of 1000 images in adverse weather conditions from real traffic environments. ACDC presents clear advantages in the total number of annotated humans and vehicles in challenging adverse conditions, thereby offering a broader spectrum of diverse scenes under challenging adverse conditions.
@@ 5.1 Domain-Adaptive Semantic Segmentation
Among the DeepLabv2-based methods, CISS, MRNet and FDA excel on the target-domain test set of ACDC-Reference, as they match (MRNet and FDA) or even exceed (CISS) the mean IoU performance of the oracle model which is trained on labeled images from ACDC-Reference. Thus, we conclude that when focusing on this earlier architecture, the domain gap which is caused by the different sensor characteristics between Cityscapes and ACDC is possible to be closed by state-of-the-art UDA methods. Considering the more recent, SegFormer-based methods, the difference in performance between the source model and the oracle model becomes only slight, namely 4.0% in mean IoU. The three top-performing methods, namely CISS, MIC and HRDA, significantly surpass the performance of the oracle model, indicating that in this domain adaptation setting, the inductive biases which are introduced to the models by the respective aforementioned domain adaptation and generalization strategies boost the target-domain performance even more than the access to in-domain training data, which only the oracle model enjoys.
@@ 7.3 Panoptic Segmentation
In Table XIX, we report the performance of Cityscapes-pre-trained panoptic segmentation models on ACDC adverse-condition images. We also observe a performance drop caused by the domain shift between Cityscapes and ACDC on state-of-the-art panoptic segmentation models. Moreover, the transformer-based Mask2Former outperforms the convolutional methods PanopticFPN and Panoptic-Deeplab. Interestingly, we also find that both mask-based methods, including K-Net and Mask2Former, present a substantially better generalization ability than per-pixel classification based methods. This indicates the mask-based methods are less affected by the domain shift in adverse conditions. However, even though the transformer-based Mask2Former model pre-trained on normal conditions shows impressive robustness to adverse conditions, it still performs much worse than the Mask2Former model which has been specifically trained on adverse conditions. By introducing the panoptically annotated extension of our ACDC dataset, we hope that the latter will contribute to closing this performance gap by fostering the development of both normal-to-adverse panoptic segmentation adaptation methods and more robust and generalizable supervised panoptic segmentation methods.
@@ 8 Uncertainty-Aware Semantic Segmentation
Existing works that model uncertainty in semantic segmentation [101, 102] are evaluated only with IoU, which does not assess the predicted confidence. In contrast, for uncertainty-aware semantic segmentation, algorithms are required to output both a hard semantic prediction and a confidence map with values in the range . The average UIoU (AUIoU) metric is computed by thresholding at multiple thresholds across the range , calculating the UIoU [9] for each threshold and averaging the results. A pixel with confidence value below the examined threshold is treated as invalid and contributes positively if (true invalid) and negatively if (false invalid).
@@ 8.1 Baselines and Oracles
We present the results of straightforward baselines for uncertainty-aware semantic segmentation that are based on methods for standard semantic segmentation in Table XX. We first evaluate three state-of-the-art methods using confidence maps that are constant and equal to 1, i.e., not modeling confidence. In this case, AUIoU reduces to IoU. Any sensible method that models confidence should improve upon this baseline. Using the max-softmax scores output by these methods as confidence maps generally yields inferior results to globally constant confidence, as softmax is not a good proxy for confidence. An upper bound for the performance of the examined methods is obtained by using a confidence oracle. More specifically, we use the binary complement of the ground-truth invalid mask as the confidence prediction. This raises AUIoU performance significantly across all conditions compared to the globally constant confidence baseline. The performance gap between the oracle and the baseline is largest for night, indicating that explicitly modeling uncertainty has the potential to improve performance especially in the nighttime domain.
@@ A.1.1 Domain adaptive semantic segmentation
For the comparison in Table 2, we use as source-domain model the DeepLabv2 [5] model that is used as the Cityscapes oracle in AdaptSegNet [43], with a performance of 65.1% mIoU on the Cityscapes validation set. For all eight unsupervised domain adaptation (UDA) methods that are compared, we use their default training configurations, including the learning rate schedule and the weights of the various losses. The number of training iterations run for each method as well as the number of self-supervised learning rounds that are used by some of the methods are reported in Table XXI. For FDA, SIM and MRNet, we run a first training round without self-training followed by a second training round with self-training, as per default implementation of these methods. For FDA, we train three separate models in each training round, one for each different value of the parameter from the set , and use the average prediction of the three models at test time. In all cases, we use the model weights corresponding to the final training iteration for testing.
The same source-domain model is also used for the experiment on adaptation to individual conditions presented in Table 3. Again, we use the default training configurations for all examined methods and across all four conditions. The number of training iterations run for each method to adapt to each condition as well as the number of self-supervised learning rounds that are used by some of the methods are reported in Table XXII. For MRNet and fog, the self-supervised training round includes 35k iterations instead of 40k. In addition, for MRNet and rain, the first training round without self-supervised training includes 25k iterations instead of 40k.
@@ A.1.2 Domain adaptive object detection
For the comparison in Table 5, we use the representative FCOS and Faster R-CNN as the source-domain models for object detection. For a fair and consistent comparison, each model is trained with a ResNet-50 backbone. For all compared UDA obejct detection methods, we use their default training configurations for Cityscapes to Foggy Cityscapes adaptation task as it is a common normal-to-adverse setting in existing UDA object detection works. All hyperparameters including the learning rate scheduling, the loss weights and the training iterations are consistent with the original configurations. Following SIGMA [33], we use the ACDC validation set for each condition to select the model weights for testing.
@@ A.2.1 Supervised Semantic Segmentation
For training the four semantic segmentation methods that are compared in Tables 9 and 10, we have generally used the default configuration for each method both in the case of condition experts and uber models. For DeepLabv2 [5], we use the architecture employed in AdaptSegNet [43] in the context of domain adaptation and not the original architecture. We have used the default learning rate schedule for each method, with the base learning rates that are reported in Table XXIII. We generally use 60 training epochs for all four methods, which yields 96k training iterations for uber models and 24k training iterations for condition experts. Exceptions to this rule are RefineNet and fog where we use 30 epochs, DeepLabv2 and fog where we use 45 epochs, DeepLabv2 and night where we use 240 epochs, and the DeepLabv3+ uber model for which we use 30 epochs. For HRNet, we use the snapshot with the best mIoU performance on the respective validation set of ACDC for predicting on the test set, while for the rest of the methods we use the final training snapshot for the same purpose.
@@ A.2.2 Supervised Instance Segmentation
For training the four instance segmentation methods that are compared in Tables 13 and 14, we have generally used the default configuration for each method both in the case of condition experts and uber models. We use the consistent ResNet-50 backbone for each model and train each model on data of each condition for 60 epoches. We use the model weights corresponding to the final training iteration for testing.
@@ A.2.3 Supervised Panoptic Segmentation
For training the four panoptic segmentation methods that are compared in Tables 15 and 16, we have generally used the default configuration for each method both in the case of condition experts and uber models. We also use the consistent ResNet-50 backbone for each model and train each model on data of each condition for 60 epoches. The model weights corresponding to the final training iteration are reported for testing.
@@ B.4 Uncertainty-aware Semantic Segmentation
In Tables LIX–LXIII, we present the class-level average uncertainty-aware IoU (AUIoU) performance of the baselines and oracles that are examined in Table XX. More specifically, Table LIX considers methods trained jointly on all conditions of ACDC and also evaluated jointly on all conditions, while Tables LX–LXIII present methods trained and evaluated on individual conditions. The results corresponding to the baseline that uses constant confidence equal to 1 are omitted, as they are identical by definition to IoU results and are thus already included in Table IX of the main paper and Tables XXXII–XXXV.
@@ C.3 Annotation
-out of which Valid | 84.93 | 96.77 | 88.16 | |||
-out of which Invalid | 7.54 | 0 | 0 | 5.48 | ||

### v5

@@ 3.5 Instance Annotation
The total 4006 adverse-condition images and 1503 normal-condition reference images are annotated at the instance level by a professional team of annotators. Each final instance-level annotation underwent a QC process. Specifically, QC was carried out through manual inspection by two individuals: the original annotator and an independent reviewer who had not previously seen the condition, in order to minimize bias. The total time required for annotating a single image given the initial semantic mask was 0.6 h on average.
@@ 3.6 Comparison to Related Datasets
To the best of our knowledge, ACDC constitutes the largest adverse-condition driving dataset for dense semantic perception to date. In Table I, we compare ACDC to prior datasets that also address semantic segmentation under adverse conditions. Most of these datasets focus on a single condition and are of small scale. WildDash covers a wider variety of adverse conditions but also has a small scale. BDD100K includes 10000 images with semantic segmentation annotations. We inspected these images manually to identify those that pertain to fog, night, rain, and snow. We found that only 1346/10000 images pertain to any of these four conditions. By contrast, ACDC is primarily composed of these four common adverse conditions. Notably, it contains one order of magnitude more annotated images than any other competing dataset for each of fog, night and rain. At the same time, our specialized annotation protocol using corresponding normal-condition images ensures reliable annotations even for invalid regions contrary to BDD100K, making ACDC a high-quality dataset for training and evaluation for adverse conditions. We also provide a comparison on the number of annotated instances with existing datasets covering adverse conditions in Table II. Only a small portion of BDD100K provides instance-level annotations in adverse conditions including fog, night, rain and snow. The DAWN dataset consists of 1000 images in adverse weather conditions from real traffic environments. ACDC presents clear advantages in the total number of annotated humans and vehicles in challenging adverse conditions compared to both DAWN and BDD100K, thereby offering a broader spectrum of diverse scenes under challenging adverse conditions.
@@ 4.2 Domain-Adaptive Object Detection
4.3 Analysis for Normal-to-Adverse Adaptation
In this section, we have investigated cross-domain semantic segmentation and object detection methods for adapting from normal conditions to adverse ones, a scenario characterized by a large domain gap. Table III shows that more recent transformer-based methods using SegFormer-type backbones [89] exhibit greater robustness than older CNN-based methods based on DeepLabv2 [66] under large domain shifts, consistently achieving higher performance. Furthermore, among various domain adaptation strategies, self-training proves beneficial across both semantic segmentation and object detection when the domain gap is large. Moreover, approaches involving input-level adaptation, such as FDA [25] and CISS [47], benefit semantic segmentation performance consistently across backbones and conditions. From a weather-specific perspective, we find that nighttime conditions present the hardest challenge across the examined tasks, whereas rain is the easiest setting for semantic segmentation and snow is the easiest one for object detection. Since object detection primarily focuses on foreground objects, our experimental results suggest that different categories experience distinct domain shifts even within the same scene. This observation highlights the potential of category-aware or instance-aware adaptation strategies to further enhance the robustness of perception models under adverse conditions.
@@ 5.1 Domain-Adaptive Semantic Segmentation
Among the DeepLabv2-based methods, CISS, MRNet and FDA excel on the target-domain test set of ACDC-Reference, as they match (MRNet and FDA) or even exceed (CISS) the mean IoU performance of the oracle model which is trained on labeled images from ACDC-Reference. Thus, we conclude that when focusing on this earlier architecture, the domain gap which is caused by the different sensor characteristics between Cityscapes and ACDC is possible to be closed by state-of-the-art UDA methods. Considering the more recent, SegFormer-based methods, the difference in performance between the source model and the oracle model becomes only slight, namely 4.0% in mean IoU. The three top-performing methods, namely CISS, MIC and HRDA, significantly surpass the performance of the oracle model, indicating that in this domain adaptation setting, the inductive biases which are introduced to the models by the respective aforementioned domain adaptation and generalization strategies boost the target-domain performance even more than access to in-domain training data does, which only the oracle model enjoys.
@@ 5.2 Domain-Adaptive Object Detection
5.3 Analysis for Sensor-Level Adaptation
This section has examined adaptation from Cityscapes to ACDC-Reference. Since both datasets contain urban images captured under clear weather, the domain gap is considerably smaller than in the normal-to-adverse adaptation setting. Consistent with the previous analysis in Sec. 4.3, self-training–based methods as well as input-level adaptation methods—two sets which are not mutually exclusive here—achieve the best performance either on semantic segmentation or object detection under a small domain gap. Interestingly, under this less complex domain shift, these types of approaches even surpass the oracle model in the semantic segmentation task, whereas domain-adaptive object detection methods fail to do so against their competing oracles. We attribute this difference to the disparity in supervision density between the two tasks: semantic segmentation provides dense pixel-level labels, while object detection offers only sparse bounding-box annotations. This finding underscores the importance of developing task-specific adaptation strategies that account for the supervision granularity of different perception tasks.
@@ 6.3 Panoptic Segmentation
6.4 Analysis for Supervised Learning on Adverse Conditions
Comparing the results from Sec. 4 and the present section, we deduce that simply leveraging direct supervision on adverse-condition data is more effective for semantic segmentation in adverse conditions than substantially upgrading the network backbone and architecture. Moreover, we have demonstrated that training a generalist semantic, instance, or panoptic segmentation model jointly on all normal- and adverse-condition data generally benefits performance across all domains, both the individual adverse-condition ones and the normal-condition one, compared to training separate condition specialists, with the exception of panoptic segmentation and rain. This observation holds despite the fact that the adverse-condition data are relatively small. Across semantic segmentation, instance segmentation, and panoptic segmentation, our results show that transformer-based methods consistently exhibit stronger generalization than CNN-based methods when trained on the same data. For CNN-based architectures, high-resolution feature representations enhance robustness across diverse domains, while effective multi-scale feature fusion is critical for strong in-domain performance. With carefully designed mechanisms, CNN-based methods can achieve performance close to that of transformer-based models in various perception tasks, narrowing the performance gap despite their inherent differences in architecture. Regarding performance on individual classes across different conditions, we have examined the semantic segmentation setting in class-specific detail in Table XI and found a non-negligible correlation between the intensity of the effect of each adversity on different classes and the respective class-level performance. Finally, we have evidenced in Table XII for the task of semantic segmentation that specialist models trained on harder conditions generalize better to different yet easier conditions than vice versa.
@@ 7.3 Panoptic Segmentation
In Table XIX, we report the performance of Cityscapes-pre-trained panoptic segmentation models on ACDC adverse-condition images. We also observe a performance drop caused by the domain shift between Cityscapes and ACDC on state-of-the-art panoptic segmentation models. Moreover, the transformer-based Mask2Former outperforms the convolutional methods PanopticFPN and Panoptic-Deeplab. Interestingly, we also find that both mask-based methods, including K-Net and Mask2Former, present a substantially better generalization ability than per-pixel classification based methods. This indicates the mask-based methods are less affected by the domain shift in adverse conditions. However, even though the transformer-based Mask2Former model pre-trained on normal conditions shows impressive robustness to adverse conditions, it still performs much worse than the Mask2Former model which has been specifically trained on adverse conditions. By introducing the panoptically annotated extension of ACDC, we hope that the latter will help to close this performance gap by fostering the development of both normal-to-adverse panoptic segmentation adaptation methods and better-generalizing supervised panoptic segmentation methods.
@@
7.4 Analysis for Externally Pre-trained Models
Consistent with the observations in Sec. 6.4, transformer-based methods demonstrate stronger generalization than CNN-based methods. In Sec. 6.4, we noted that certain CNN-based models can achieve substantial in-domain performance gains through carefully designed feature aggregation mechanisms. However, in the present externally pre-trained setting, such designs may reduce out-of-distribution robustness. For example, HTC underperforms its base model, Cascade Mask R-CNN, in Table XVIII, and Panoptic-DeepLab falls behind its base PanopticFPN in Table XIX on out-of-distribution evaluations.
@@ 8 Uncertainty-Aware Semantic Segmentation
Existing uncertainty-based works on semantic segmentation [102, 103] are evaluated only with IoU, which does not assess the predicted confidence. In contrast, for uncertainty-aware semantic segmentation, algorithms are required to output both a hard semantic prediction and a confidence map with values in the range . The average UIoU (AUIoU) metric is computed by thresholding at multiple values across the range , calculating the UIoU [9] for each threshold and averaging the results. A pixel with confidence value below the threshold at hand is treated as invalid and contributes positively if (true invalid) and negatively if (false invalid).
@@ 8.1 Baselines and Oracles
We present the results of straightforward baselines for uncertainty-aware semantic segmentation that are based on methods for standard semantic segmentation in Table XX. We first evaluate three such methods using constant confidence maps equal to 1, not modeling confidence. In this case, AUIoU reduces to IoU. Any sensible method that models confidence should improve upon this baseline. Using the max-softmax scores of these methods as confidence maps generally yields inferior results to globally constant confidence, as softmax is not a good proxy for confidence. An upper bound for the performance of the examined methods is obtained by using a confidence oracle. More specifically, we use the binary complement of the ground-truth invalid mask as the confidence prediction. This raises AUIoU performance significantly across all conditions compared to the globally constant confidence baseline. The performance gap between the oracle and the baseline is largest for night, indicating that explicitly modeling uncertainty has the potential to improve performance especially in the nighttime domain. We have also trained [102] on ACDC, using the GT invalid masks for training its outlier detection part. The learned confidence by [102] leads to lower test set AUIoU (52.0%) than constant confidence (53.0%), indicating that a better modeling of uncertainty is needed in future approaches.
@@ A.1.1 Domain adaptive semantic segmentation
For the comparison in Table III, we use as source-domain model the DeepLabv2 [66] model that is used as the Cityscapes oracle in AdaptSegNet [16], with a performance of 65.1% mIoU on the Cityscapes validation set. For all eight unsupervised domain adaptation (UDA) methods that are compared, we use their default training configurations, including the learning rate schedule and the weights of the various losses. The number of training iterations run for each method as well as the number of self-supervised learning rounds that are used by some of the methods are reported in Table XXI. For FDA, SIM and MRNet, we run a first training round without self-training followed by a second training round with self-training, as per default implementation of these methods. For FDA, we train three separate models in each training round, one for each different value of the parameter from the set , and use the average prediction of the three models at test time. In all cases, we use the model weights corresponding to the final training iteration for testing.
The same source-domain model is also used for the experiment on adaptation to individual conditions presented in Table IV. Again, we use the default training configurations for all examined methods and across all four conditions. The number of training iterations run for each method to adapt to each condition as well as the number of self-supervised learning rounds that are used by some of the methods are reported in Table XXII. For MRNet and fog, the self-supervised training round includes 35k iterations instead of 40k. In addition, for MRNet and rain, the first training round without self-supervised training includes 25k iterations instead of 40k.
@@ A.1.2 Domain adaptive object detection
For the comparison in Table V, we use the representative FCOS and Faster R-CNN as the source-domain models for object detection. For a fair and consistent comparison, each model is trained with a ResNet-50 backbone. For all compared UDA object detection methods, we use their default training configurations for Cityscapes to Foggy Cityscapes adaptation task as it is a common normal-to-adverse setting in existing UDA object detection works. All hyperparameters including the learning rate scheduling, the loss weights and the training iterations are consistent with the original configurations. Following SIGMA [34], we use the ACDC validation set for each condition to select the model weights for testing.
@@ A.2.1 Supervised Semantic Segmentation
For training the four semantic segmentation methods that are compared in Tables 9IX and X, we have generally used the default configuration for each method both in the case of condition experts and uber models. For DeepLabv2 [66], we use the architecture employed in AdaptSegNet [16] in the context of domain adaptation and not the original architecture. We have used the default learning rate schedule for each method, with the base learning rates that are reported in Table XXIII. We generally use 60 training epochs for all four methods, which yields 96k training iterations for uber models and 24k training iterations for condition experts. Exceptions to this rule are RefineNet and fog where we use 30 epochs, DeepLabv2 and fog where we use 45 epochs, DeepLabv2 and night where we use 240 epochs, and the DeepLabv3+ uber model for which we use 30 epochs. For HRNet, we use the snapshot with the best mIoU performance on the respective validation set of ACDC for predicting on the test set, while for the rest of the methods we use the final training snapshot for the same purpose.
@@ A.2.2 Supervised Instance Segmentation
For training the four instance segmentation methods that are compared in Tables XIII and XIV, we have generally used the default configuration for each method both in the case of condition experts and uber models. We use the consistent ResNet-50 backbone for each model and train each model on data of each condition for 60 epochs. We use the model weights corresponding to the final training iteration for testing.
@@ A.2.3 Supervised Panoptic Segmentation
For training the four panoptic segmentation methods that are compared in Tables XV and XVI, we have generally used the default configuration for each method both in the case of condition experts and uber models. We also use the consistent ResNet-50 backbone for each model and train each model on data of each condition for 60 epochs. The model weights corresponding to the final training iteration are reported for testing.
@@ B.4 Uncertainty-aware Semantic Segmentation
In Tables LIX–LXIII, we present the class-level average uncertainty-aware IoU (AUIoU) performance of the baselines and oracles that are examined in Table XX. More specifically, Table LIX considers methods trained jointly on all conditions of ACDC and also evaluated jointly on all conditions, while Tables LX–LXIII present methods trained and evaluated on individual conditions. The results corresponding to the baseline that uses constant confidence equal to 1 are omitted, as they are identical by definition to IoU results and are thus already included in Table IX and Tables XXXII–XXXV.
@@ C.3 Annotation
| -out of which Valid | 84.93 | 96.77 | 88.16 | |||
| -out of which Invalid | 7.54 | 0 | 0 | 5.48 | ||